OpenStudy (jekdidkdjrjrjjr):

What is the correct classification of the system of equations below? 12x − 3y = 9 y − 4x = -3 A) parallel B) coincident C) intersecting

OpenStudy (jekdidkdjrjrjjr):

@fluttershyK @rebeccaxhawaii @thomas5267 @AloneS @xGuardians @ijal @SamandDean1967 @Atrineas @latinC @arianna1453 @linn99123 @calculusxy @welshfella @MathMusician @marziman @supersmart1001 @studio-games @Erak @taylarm7 @ijal @mathstudent55 @Vocaloid @fluttershyK @sweetburger @narissa @HaileyD @thinkerblinker @YumYum247 @pigsruleyeah @Testing123guest @shaniehh @Stbest @FortyTheRapper @Mehek14 @volleyballlover55 @heacain @robtobey @IHelpYouLearn @Lilly_Jarzen @dmezzullo

OpenStudy (anonymous):

coincident

OpenStudy (jekdidkdjrjrjjr):

lmao which one is it @heacain

OpenStudy (jekdidkdjrjrjjr):

2 ppl saying different answers idk

OpenStudy (erak):

Yeah it's coincident because they would give the same equation. 12x - 3y = 9 -3y = -12x + 9 y = 4x - 3

OpenStudy (anonymous):

it would be C because they line up exactly on a graph

OpenStudy (erak):

It would be B. It doesn't intersect, it's the same line.

OpenStudy (latinc):

between B,A, or C it has to be one of this